# SDDBench: A Benchmark for Synthesizable Drug Design

### v1

@@ Abstract
A significant challenge in wet lab experiments with current drug design generative models is the trade-off between pharmacological properties and synthesizability. Molecules predicted to have highly desirable properties are often difficult to synthesize, while those that are easily synthesizable tend to exhibit less favorable properties. As a result, evaluating the synthesizability of molecules in general drug design scenarios remains a significant challenge in the field of drug discovery. The commonly used synthetic accessibility (SA) score aims to evaluate the ease of synthesizing generated molecules, but it falls short of guaranteeing that synthetic routes can actually be found. Inspired by recent advances in top-down synthetic route generation, we propose a new, data-driven metric to evaluate molecule synthesizability. Our approach directly assesses the feasibility of synthetic routes for a given molecule through our proposed round-trip score. This novel metric leverages the synergistic duality between retrosynthetic planners and reaction predictors, both of which are trained on extensive reaction datasets. To demonstrate the efficacy of our method, we conduct a comprehensive evaluation of round-trip scores alongside search success rate across a range of representative molecule generative models. Code is available at https://github.com/SongtaoLiu0823/SDDBench.
@@ 1 Introduction
Drug design is a fundamental problem in machine learning for drug discovery. However, when these computationally predicted molecules are put to the test in wet lab experiments, a critical issue often arises: many of them prove to be unsynthesizable in practice. This synthesis gap can be attributed to two primary factors. Firstly, while structurally feasible, the predicted molecules often lie far beyond the known synthetically-accessible chemical space. This significant departure from known chemical territory makes it extremely difficult, and often impossible, to discover feasible synthetic routes. This synthesis challenge is underscored by numerous clinical drugs derived from natural products, which, due to their intricate structures, can only be obtained through direct extraction from natural sources rather than synthesis methods (Zheng et al., 2022). These natural products often have complex ring structures and multiple chiral centers, which makes their chemical synthesis challenging. Additionally, the biological processes that create these compounds are frequently not well understood, increasing the complexity of laboratory synthesis. Secondly, even when plausible reactions are identified based on literature, they may fail in practice due to the inherent complexity of chemistry. The sensitivity of chemical reactions is such that even minor change in functional groups can potentially prevent a reaction from happening as anticipated.
@@
The ability to synthesize designed molecules on a large scale is crucial for drug development. Current methods (You et al., 2018; Gao & Coley, 2020) rely on the Synthetic Accessibility (SA) score (Ertl & Schuffenhauer, 2009) for synthesizability evaluation. This score assesses how easily a drug can be synthesized by combining fragment contributions with a complexity penalty. However, this metric has limitations as it evaluates synthesizability based on structural features and fails to account for the practical challenges involved in developing actual synthetic routes for these molecules. Therefore, we propose that a more comprehensive gold standard for assessing synthesizability (Stanley & Segler, 2023; Hastedt et al., 2024) should be the ability to identify a feasible synthetic route for each generated molecule.
@@
In this work, we introduce a novel metric inspired by recent advancements in retrosynthesis. Our metric evaluates the ease of predicting synthetic routes for molecules using data-driven retrosynthetic planners trained on existing reaction data. Unlike the approaches (Bradshaw et al., 2019b; Gao et al., 2022a), which integrate drug design and synthetic planning, we adopt a two-step process. First, we generate molecules using drug design generative models. We then employ a retrosynthetic planner to predict synthetic routes for these generated molecules. To assess the feasibility of these routes, we utilize a reaction prediction model as a simulation agent, replacing wet lab experiments. This model attempts to reproduce both the synthetic route and the generated molecule, starting from the predicted route’s starting materials. Finally, we introduce the round-trip score, which computes the Tanimoto similarity between the reproduced molecule and the originally generated molecule.
With the round-trip score as foundation, we develop a new benchmark and dataset to evaluate the “synthesizability” of molecules predicted by current structure-based drug design (SBDD) generative models. Our approach draws inspiration from evaluation methods used in image generation, such as the CLIP score (Radford et al., 2021; Hessel et al., 2021). In image generation, the CLIP score assesses the similarity between generated images and their corresponding text descriptions using the pre-trained CLIP model (Radford et al., 2021). Analogously, our point-wise round-trip score evaluates whether the starting materials in a predicted synthetic route can successfully undergo a series of reactions to produce the generated molecule. Through our analysis, we have found a significant correlation: molecules with feasible synthetic routes, as predicted by our retrosynthetic planner, consistently achieve higher round-trip scores compared to those lacking feasible routes. This finding underscores the effectiveness of our metric in assessing the practical synthesizability of computationally generated molecules. Our contributions can be summarized as follows:
We redefine molecule synthesizability from a data-centric perspective. In our view, a molecule is synthesizable if retrosynthetic planners, trained on existing reaction data, can predict a feasible synthetic route for it.
Our work is the firs to bridge the gap between drug design and retrosynthetic planning. We introduce the round-trip score, integrating retrosynthesis prediction, reaction prediction, and drug design into a unified framework. In this framework, drug design models generate new drug molecules, retrosynthetic planners predict their synthetic routes, and reaction prediction models reproduce both the predicted routes and the generated molecules.
@@ 2 Background
In this section, we discuss the details of drug design and molecule synthesis. Machine learning algorithms for molecule synthesis can be categorized into three main types: forward reaction prediction models, backward retrosynthesis prediction models, and search algorithms.
@@ 2.1 Structure-Based Drug Design
While our newly developed benchmark is capable of evaluating a wide range of drug design models, this work specifically focuses on assessing the synthesizability of molecules generated by SBDD models. The primary goal of SBDD is to generate ligand molecules capable of binding to a specific protein binding site. In this context, we represent the target protein and ligand molecule as and , respectively. Here, and denote the number of atoms in the protein and ligand . For each atom, represents its position in three-dimensional space, while encodes its type. The core challenge of SBDD lies in accurately modeling the conditional distribution , which describes the likelihood of a ligand molecule given a specific protein structure.
@@ 2.3 Retrosynthesis Prediction
Retrosynthesis, the inverse process of reaction prediction, aims to identify a set of reactants capable of synthesizing a given product molecule through a single chemical reaction. This process essentially works backwards from the desired product, determining the precursor molecules necessary for its synthesis. By doing so, retrosynthesis plays a crucial role in planning synthetic routes for complex molecules, particularly in drug discovery and materials science.
@@
Reaction prediction and retrosynthesis prediction differ fundamentally in their nature and objectives. Reaction prediction is largely a deterministic task, where specific reactants under given conditions typically yield a predictable outcome. In contrast, retrosynthesis prediction is inherently a one-to-many task, providing multiple potential routes to a desired product as illustrated in Figure 2. To be accurate, we also use examples from high school chemistry textbooks that demonstrate how a single set of reactants can produce different products under varying conditions. An example is the reaction between hydrogen bromide (HBr) and 1,3-butadiene. As shown in Figure 1, this reaction can result in either 1,2-addition or 1,4-addition products, with the outcome determined by the specific reaction conditions employed. Note that, however, that under any given set of conditions, only one of these products will form. Retrosynthesis, on the other hand, typically does not specify reaction conditions in advance. Its goal is to identify any set of reactants capable of synthesizing the target product, considering all valid solutions. This approach offers flexibility in planning synthesis strategies.
@@ 2.5 Retrosynthetic Planning
Retrosynthetic planning is a strategic approach to predicting synthetic routes for target molecules. This process works backwards from the desired target, identifying potential precursor molecules that could be transformed into the target through chemical reactions. These precursors are then further decomposed into simpler, readily available starting materials or building blocks. A synthetic route can be formally represented as a tuple with four elements: , where is the target molecule, represents the space of starting materials, denotes the specific starting materials used, is the series of reactions leading to , and represents the intermediates. In organic synthesis, a “route” refers to the complete flowchart of reactions required to synthesize a target molecule, as illustrated in Figure 3. This definition differs from its usage in computer science. Synthetic routes can be classified as convergent (Figure 3) or non-convergent, depending on whether the reactions within the route have branching points (Gao et al., 2022a). The planning process is iterative. At each step, single-step retrosynthesis models predict various sets of potential reactants that could lead to the desired product. A search algorithm then selects the most promising solutions to extend the synthetic route further. This process continues until all leaf nodes correspond to readily available starting materials, resulting in a complete synthetic route from purchasable molecules to the target compound.
@@ 2.6 Evaluation of Molecule Synthesis
Current evaluation methods for single-step reaction and retrosynthesis predictions primarily rely on the exact match metric. This approach assesses whether the predicted results match with the ground truth in the test dataset. Typically, multiple predictions are generated, and the top-k test accuracy is reported. For single-step retrosynthesis, an additional metric known as round-trip accuracy (Schwaller et al., 2020) is often employed (Mo et al., 2021; Maziarz et al., 2023; Skoraczyński et al., 2023; Torren-Peraire et al., 2024). This method involves using the predicted reactants as input for a reaction model to generate a product, which is then compared to the original product used as input for the retrosynthesis model. In retrosynthetic planning, two primary evaluation metrics are commonly used:
Search Success Rate: This metric quantifies the proportion of molecules for which the shortest possible synthetic route, with all leaf nodes as starting materials, can be identified within a limited number of retrosynthesis iterations.
Set-wise Exact Match: This metric (Liu et al., 2023b, 2024a) addresses a key limitation of the search success rate. It verifies whether the predicted starting materials can indeed be synthesized into the target molecule through a series of reactions. This is achieved by comparing the predicted starting materials with those in reference synthetic routes extracted from reaction databases.
Researchers can more accurately assess the performance of machine learning models in retrosynthetic planning by utilizing the set-wise exact match metric. This approach helps ensure that the proposed synthetic routes are practically feasible and viable in real-world applications.
@@ 3 Round-trip Score
In this section, we introduce a novel metric called the round-trip score. This metric is designed to assess the feasibility of synthetic routes for molecules generated by drug design models. Specifically, it evaluates the probability that retrosynthetic planners, trained on current reaction data, can successfully predict feasible synthetic routes for these proposed molecules. This approach bridges the gap between drug design and practical synthesis, providing a crucial link between computational predictions and real-world applicability in drug discovery.
3.1 Problem Formulation
Our objective is to leverage generative models to predict molecules that not only satisfy a given condition , but also have synthetic routes that can be predicted by retrosynthetic planners trained on existing reaction data. This dual goal can be formally expressed as:
To our knowledge, this work represents the first attempt to bridge the gap between drug design and retrosynthetic planning. We believe that integrating these two domains into a unified problem, as shown in Eq. (1), is crucial for designing synthesizable drugs. While previous methods (Bradshaw et al., 2019b; Gao et al., 2022a) have attempted to simultaneously generate molecules and their synthetic routes, they face limitations. For instance, the approach in (Gao et al., 2022a) is constrained by the use of a limited number of reaction templates, which hinders scalability. In contrast, our approach to Eq. (1) is decoupled. We first propose a molecule using a generative model. Then, we predict its synthetic route using a retrosynthetic planner. This strategy allows us to assess the probability that the feasible synthetic routes of molecules generated by these models can be predicted by retrosynthetic planners trained on existing reaction data.
3.2 Proposed Metric: Round-trip Score
In this section, we introduce a novel metric called the round-trip score and provide a comprehensive evaluation of its reliability. This metric is designed to bridge the gap between molecular design and practical synthesis, offering a quantitative measure of a molecule’s synthesizability based on current reaction knowledge.
3.2.1 Round-trip Score
First, we explain how to design a metric that distinguishes between molecules whose feasible synthetic routes are easily predicted by retrosynthetic planners and those whose routes are not. Theoretically, this metric should be higher for molecules whose feasible synthetic routes are more readily predictable, compared to those with feasible routes that are more difficult to predict. Let measure the likelihood that the retrosynthetic planner can successfully predict the feasible synthetic route for . Therefore, for two molecules and , if , then the feasible synthetic route of is easier to predict than that of . We define the gap between the metric scores of these two molecules as follows:
where and are the model parameters which will be elaborated later. The effectiveness of the metric can be gauged by the magnitude of the gap . A larger should emerge when there’s a notable difference in how easily the retrosynthetic planner can predict feasible synthetic routes for molecules and . Essentially, a more substantial indicates that the metric successfully differentiates between distinct molecules. To quantify , we must first evaluate the feasibility of the predicted synthetic routes—determining whether the proposed route can actually synthesize the target molecule. Current approaches for assessing predicted routes in retrosynthetic planning primarily fall into two categories: the search success rate (Segler et al., 2018), and the set-wise exact match metric (Liu et al., 2023b).
Our previous discussions reveal that the search success rate is not a reliable metric. While the set-wise exact match is viable when a reference synthetic route is available, it presents a significant challenge: for newly generated molecules, reference routes are often absent from existing reaction datasets, making it difficult to evaluate the feasibility of routes predicted by retrosynthetic planners. This raises a crucial question: Can we devise a metric to assess the feasibility of predicted routes by retrosynthetic planners for molecules proposed by drug design generative models?
@@ 3.2.1 Round-trip Score
Inspired by the round-trip accuracy metric used in single-step retrosynthesis evaluation, we propose the round-trip score. Given a molecule proposed by a generative model, we first use a retrosynthetic planner to predict a synthetic route. Starting from the initial materials of this route, we then employ a reaction model to simulate wet lab experiments and reproduce the synthetic route until we reach the final molecule . Finally, we compute the Tanimoto similarity between and , which we define as the round-trip score. Figure 4 provides an illustration of the entire process. The round-trip score, which encapsulates this process, can be mathematically expressed as follows:
@@
Motivation.
The proposed metric is motivated by two key factors. Firstly, it draws inspiration from CREBM (Liu et al., 2024a), which employs a similar approach to evaluate the feasibility of starting material sets and creates a synthetic preference dataset for training an adapter, ultimately enhancing the performance of retrosynthetic planners. Secondly, there is currently a dearth of alternative methods for assessing the feasibility of predicted routes without relying on references. The only available option is to simulate the wet lab process using a reaction model. Consequently, the robustness of this metric is inherently tied to the accuracy of the underlying reaction model.
Discussion.
In this work, we primarily focus on using retrosynthetic planners to predict synthetic routes for generated molecules. It’s important to note that while some natural products used in clinical settings can theoretically be synthesized, current chemical process technologies often fall short of achieving this in practical wet lab conditions (Zheng et al., 2022). Furthermore, even for synthesizable molecules, retrosynthetic planners trained on existing reaction data may fail to predict viable synthetic routes. We attribute this limitation largely to the scarcity and proprietary nature of comprehensive reaction data. This challenge parallels similar issues in Large Language Model (LLM) research, where advancements towards Artificial General Intelligence (AGI) have been achieved through models like GPT (Radford et al., 2018) by training on increasingly larger datasets. Drawing from this parallel, we think that access to more extensive and diverse reaction data would significantly enhance our ability to design synthesizable drugs and successfully bring them to mass production. This underscores the critical importance of expanding and enriching our reaction databases to bridge the gap between theoretical possibility and practical synthesis.
3.2.2 Evaluation of the Reliability of our Round-trip Score

### v2

# Evaluating Molecule Synthesizability via Retrosynthetic Planning and Reaction Prediction

@@ Abstract
A significant challenge in wet lab experiments with current drug design generative models is the trade-off between pharmacological properties and synthesizability. Molecules predicted to have highly desirable properties are often difficult to synthesize, while those that are easily synthesizable tend to exhibit less favorable properties. As a result, evaluating the synthesizability of molecules in general drug design scenarios remains a significant challenge in the field of drug discovery. The commonly used synthetic accessibility (SA) score aims to evaluate the ease of synthesizing generated molecules, but it falls short of guaranteeing that synthetic routes can actually be found. Inspired by recent advances in top-down synthetic route generation and forward reaction prediction, we propose a new, data-driven metric to evaluate molecule synthesizability. This novel metric leverages the synergistic duality between retrosynthetic planners and reaction predictors, both of which are trained on extensive reaction datasets. To demonstrate the efficacy of our metric, we conduct a comprehensive evaluation of round-trip scores across a range of representative molecule generative models.
@@ 1 Introduction
Drug design is a fundamental problem in machine learning for drug discovery. However, when these computationally predicted molecules are put to the test in wet lab experiments, a critical issue often arises: many of them prove to be unsynthesizable in practice (Parrot et al., 2023). This synthesis gap can be attributed to two primary factors. Firstly, while structurally feasible, the predicted molecules often lie far beyond the known synthetically-accessible chemical space (Ertl & Schuffenhauer, 2009). This significant departure from known chemical territory makes it extremely difficult, and often impossible, to discover feasible synthetic routes (Segler et al., 2018; Liu et al., 2023b). This synthesis challenge is underscored by numerous clinical drugs derived from natural products, which, due to their intricate structures, can only be obtained through direct extraction from natural sources rather than synthesis methods (Zheng et al., 2022). These natural products often have complex ring structures and multiple chiral centers, which makes their chemical synthesis challenging (Paterson & Anderson, 2005). Additionally, the biological processes that create these compounds are frequently not well understood, increasing the complexity of laboratory synthesis. Secondly, even when plausible reactions are identified based on literature, they may fail in practice due to the inherent complexity of chemistry (Lipinski, 2004). The sensitivity of chemical reactions is such that even minor changes in functional groups can potentially prevent a reaction from happening as anticipated.
@@
The ability to synthesize designed molecules on a large scale is crucial for drug development. Some current methods (You et al., 2018; Gao & Coley, 2020) rely on the Synthetic Accessibility (SA) score (Ertl & Schuffenhauer, 2009) for synthesizability evaluation. This score assesses how easily a drug can be synthesized by combining fragment contributions with a complexity penalty. However, this metric has limitations as it evaluates synthesizability based on structural features and fails to account for the practical challenges involved in developing actual synthetic routes for these molecules. In other words, a high SA score does not guarantee that a feasible synthetic route for the molecule can be identified using available molecule synthesis tools (Genheden et al., 2020; Tripp et al., 2022).
To overcome the limitations of the SA score, recent works (Guo & Schwaller, 2024; Cretu et al., 2024) have employed retrosynthetic planners or AiZynthFinder (Genheden et al., 2020) to evaluate the synthesizability of generated molecules. These tools are used to find synthetic routes and assess the proportion of molecules for which routes can be found. As a result, these works rely on the search success rate for evaluating molecule synthesizability. However, this metric is overly lenient, as it fails to ensure that the proposed routes are actually capable of synthesizing the target molecules (Liu et al., 2023b). In practice, many reactions predicted by these tools may not be simulated in the wet lab, as these tools often rely on data-driven retrosynthesis models prone to predicting unrealistic or hallucinated reactions (Zhong et al., 2023; Tripp et al., 2024).
To address the overly lenient evaluation metrics in previous retrosynthesis studies, where success is often defined merely by finding a “solution” without any regard to whether the solution can be executed in the wet lab (Tripp et al., 2024), FusionRetro (Liu et al., 2023b) proposes assessing whether the starting materials111Starting materials are defined as commercially purchasable molecules. ZINC (Sterling & Irwin, 2015) provides open-source databases of purchasable compounds, and we define the compounds listed in these databases as our starting materials. of a predicted route of a target molecule match those in reference routes from the literature database for a target molecule. However, for new molecules generated by drug design models, reference synthetic routes are often unavailable in literature databases. This raises a critical question:
Can data-driven retrosynthetic planners be used to evaluate the synthesizability of these molecules?
Inspired by recent advancements that leverage forward reaction models (Sun et al., 2021) to enhance retrosynthesis algorithms and rank the top-k synthetic routes predicted by retrosynthetic planners (Schwaller et al., 2019b; Liu et al., 2024), we propose a three-stage approach that incorporates forward reaction models for evaluating molecule synthesizability to address this question.
@@
Our evaluation process consists of three stages. In the first stage, we use a retrosynthetic planner to predict synthetic routes for molecules generated by drug design generative models. In the second stage, we assess the feasibility of these routes using a reaction prediction model as a simulation agent, serving as a substitute for wet lab experiments. This model attempts to reconstruct both the synthetic route and the generated molecule, starting from the predicted route’s starting materials. In the third stage, we calculate the Tanimoto similarity, also called the round-trip score, between the reproduced molecule and the originally generated molecule as the synthesizability evaluation metric. Our point-wise round-trip score evaluates whether the starting materials can successfully undergo a series of reactions to produce the generated molecule.
With the round-trip score as the foundation, we develop a new benchmark to evaluate the “synthesizability” of molecules predicted by current structure-based drug design (SBDD) generative models. Our contributions can be summarized as follows:
We recognize the limitations of the current metrics used for evaluating molecule synthesizability. Therefore, we propose the round-trip score as a metric to evaluate the synthesizability of new molecules generated by drug design models.
@@ 2.1 Structure-Based Drug Design
While our newly developed benchmark is capable of evaluating a wide range of drug design models, this work specifically focuses on assessing the synthesizability of molecules generated by SBDD models. The primary goal of SBDD is to generate ligand molecules capable of binding to a specific protein binding site.
@@ 2.3 Retrosynthesis Prediction
Retrosynthesis aims to identify a set of reactants capable of synthesizing a given product molecule through a single chemical reaction. This process essentially works backward from the desired product, determining the precursor molecules necessary for its synthesis. By doing so, retrosynthesis plays a crucial role in planning synthetic routes for complex molecules, particularly in drug discovery and materials science.
@@
Reaction prediction and retrosynthesis prediction differ fundamentally in their nature and objectives. Reaction prediction is a deterministic task, where specific reactants under given conditions typically yield a predictable outcome. In contrast, retrosynthesis prediction is inherently a one-to-many task, providing multiple potential routes to a desired product as illustrated in Figure 1.
@@ 2.5 Retrosynthetic Planning
Retrosynthetic planning aims to predict synthetic routes for target molecules. This process works backward from the desired target, predicting potential precursor molecules that could be transformed into the target through chemical reactions. These precursors are then further decomposed into simpler, readily available starting materials. A synthetic route can be formally represented as a tuple with four elements: , where is the target molecule, represents the space of starting materials, denotes the specific starting materials used, is the series of reactions leading to , and represents the intermediates. In organic synthesis, a “route” refers to the complete flowchart of reactions required to synthesize a target molecule, as illustrated in Figure 1. This definition differs from its usage in computer science. Synthetic routes can be classified as convergent (Figure 1) or non-convergent, depending on whether the reactions within the route have branching points (Gao et al., 2022). The planning process is iterative. At each step, single-step retrosynthesis models predict various sets of potential reactants that could lead to the desired product. A search algorithm then selects the most promising solutions to extend the synthetic route further. This process continues until all leaf nodes correspond to readily available starting materials, resulting in a complete synthetic route from purchasable molecules to the target compound.
@@ 2.6 Evaluation of Molecule Synthesis
Current evaluation methods for single-step reaction and retrosynthesis predictions rely on the exact match metric. This approach assesses whether the predicted results match the ground truth in the test dataset. Multiple predictions are generated, and the top-k test accuracy is reported.
Until recently, evaluation criteria for retrosynthetic planning had not reached a clear consensus, but they have now converged on a few key metrics. Among these, one of the most widely used is the success rate of finding a synthetic route within a limited number of calls for single-step retrosynthesis prediction (typically capped at 500). However, this metric, known as the search success rate, is overly lenient as it does not verify whether the searched synthetic route can be executed in the wet lab to synthesize the target molecule. This limitation is particularly problematic for targets requiring long synthetic routes, where errors can accumulate across multiple steps.
To illustrate this limitation, we observe that existing single-step models achieve top-5 accuracies of less than 80% (Somnath et al., 2021). In contrast, current retrosynthetic planning methods report search success rates exceeding 99% (Xie et al., 2022) under the limit of 500 single-step retrosynthesis prediction iterations. This discrepancy is counterintuitive since longer synthetic routes should inherently have a lower likelihood of success due to the increasing complexity of the synthesis. This raises concerns about the quality of the routes deemed “successful” by multi-step planners.
To address the limitations of using search success rate as an evaluation metric for retrosynthetic planning, FusionRetro (Liu et al., 2023b) introduces a matching-based evaluation approach. This method compares the starting materials of synthetic routes predicted by retrosynthetic planners for target molecules with those from reference routes retrieved from literature databases. If the starting materials of a predicted route match those of any reference route, the prediction is considered accurate and successful. This approach aligns with the evaluation methodology used in single-step retrosynthesis, which also relies on literature databases.
Moreover, FusionRetro goes further by constructing a reaction network from all reactions in the literature database. For a given target molecule, it extracts all synthetic routes from this reaction network, with the leaf nodes within these routes being the starting materials. As FusionRetro reports, this often enables the identification of multiple synthetic routes for a target molecule within the literature database.
While this matching-based approach has its limitations such as the inability of current literature databases to cover all equivalent synthetic routes to a target molecule, this limitation is not unique to FusionRetro and also applies to existing retrosynthesis evaluation methods. Despite this, the matching-based evaluation metric is a more reliable and rigorous alternative to the search success rate.
To illustrate the differences between the two metrics, we introduce an example in Figure 2. The top part of the figure shows a predicted synthetic route where the retrosynthesis prediction within the highlighted circle is incorrect, while the bottom part illustrates a correctly predicted synthetic route. Despite this, both routes are considered successful under the search success rate metric. This is because the search success rate evaluates the solvability of finding a route with leaf nodes as starting materials within a limited number of single-step retrosynthesis prediction iterations for a given target molecule.
In contrast, the starting material matching-based metric clearly distinguishes between the two routes. It identifies the top route as incorrect and the bottom route as correct, as the incorrect route in the top example would not match any entries in the literature reaction database. This example intuitively demonstrates that the matching-based evaluation metric provides a more reliable and accurate assessment than the search success rate.
@@ 3 Round-trip Score
In this section, we introduce a novel metric called the round-trip score. This metric is designed to assess the feasibility of synthetic routes for molecules generated by drug design models.
3.1 Motivation
As discussed in Sections 1 and 2, current heuristics-based metrics for evaluating molecule synthesizability, such as the SA score, fail to ensure that synthetic routes can be identified using existing data-driven molecule synthesis tools. However, these tools typically rely on the search success rate as a metric, which does not assess the feasibility of the predicted routes. We have identified a critical flaw in the search success rate metric in Section 2.6 and find that the match-based evaluation metric provides a more reliable alternative.
However, for new molecules generated by drug design generative models, reference routes are often missing from literature databases, making it impossible to evaluate predicted routes using match-based metrics. Ideally, the most accurate evaluation would involve directly validating the predicted routes in the wet lab to confirm whether they can synthesize the target molecules. However, this approach is prohibitively expensive, especially when evaluating large numbers of molecules.
To address this challenge, we note that recent forward reaction prediction models achieve top-1 accuracies exceeding 90% (Bi et al., 2021). These models can simulate reactions to verify whether the predicted synthetic routes are capable of synthesizing the target molecules. While this approach has its limitations, it is far more reliable than the search success rate and avoids the significant costs associated with wet lab experiments.
3.2 Three-stage Evaluation Process
@@
Given a molecule proposed by a generative model, we first use a retrosynthetic planner to predict a synthetic route. Starting from the initial materials of this route, we then employ a reaction model to simulate wet lab experiments and reproduce the synthetic route until we reach the final molecule . Finally, we compute the Tanimoto similarity between and , which we define as the round-trip score. Figure 3 provides an illustration of the entire process. The round-trip score, which encapsulates this process, can be mathematically expressed as follows:
@@
4 Experiments
Our experiment consists of two parts. The first part focuses on assessing the reliability of the SA score, search success rate and round-trip score. The second part evaluates the synthesizability of generated molecules using the round-trip score.
4.1 Evaluating the Reliability of Synthesizability Metrics
Currently, retrosynthetic planners are employed to generate synthetic routes for new molecules. When the planner predicts a route, our synthesizability evaluation metrics need to differentiate between feasible and infeasible routes. Therefore, we need a dataset to assess such discriminative capability of these synthesizability evaluation metrics.
Dataset Construction.
To prepare the dataset, we first clean and deduplicate the USPTO reactions, resulting in approximately 916k reactions. These reactions are then used to construct a reaction network. Molecules with an out-degree of 0 in the network are treated as target molecules, and their corresponding synthetic routes are extracted. This process yields synthetic routes for 107,354 molecules, where the leaf nodes in the routes are starting materials. Note that some molecules can be synthesized through multiple synthetic routes in the dataset.
The dataset is divided into training, validation, and test sets with a ratio of 98%, 1%, and 1%, respectively, based on the molecules. These splits consist of 105,218, 1,068, and 1,068 data points, respectively. Each data point includes the target molecule and all its associated synthetic routes.